# Time-Series Foundation Model for Value-at-Risk Forecasting

### v6

@@ 5.1 Actual over Expected Ratio
Table 2 reports the summary statistics over the out-of-sample period from January 2015 to September 202, across 92 stocks, of the $|1-\text{AE}|$ values for the VaR forecasts at different levels, and the corresponding models in the out-sample period. Specifically, we report the minimum, mean, median, maximum, and standard deviation across all the stocks. Additionally, we report the count of stocks for which each of the considered models was the best (achieved lowest value of $|1-AE|$) or was within top two models (1st-2nd best). Note that, as mentioned before, the VaR forecasts for TimesFM pretrained model are only available at $10\%$, there we have three extra models in the last block of Table 2.
@@
Tables 4, 5 and 6 present the out-of-sample results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The out-of-sample results are also depicted via bar plots in Figures 1-3. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. The higher the count, the better the model. From these tables, it is evident that the TimesFM shows promising results; in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or performs fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lengths show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@ 6 Conclusion and discussion
Determining whether pre-trained foundation time-series models can outperform traditional econometric methods is a key question, as their success would mean that anyone with minimal statistical and mathematical knowledge could achieve relatively good VaR forecasts. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training that yielded suboptimal results. However, fine-tuning the model with domain-specific data led to significant statistical improvements: In terms of Actual vs. Expected violations (AE), fine-tuned TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in the recent econometric literature (Patton et al., 2019). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as foundation models can be utilized with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al., 2024). This lack of transparency is especially problematic in high-stakes financial settings, where understanding the reasoning behind predictions is essential for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) forecasting by offering a completely new alternative to traditional econometric methods. We anticipate future research addressing various econometric and financial market challenges, such as portfolio optimization and volatility forecasting, as well as applications in other fields where time series forecasting methods are required. These include forecasting electricity and commodity prices Berrisch and Ziel (2024); Hasselgren et al. (2025), macroeconomic variables Čapek et al. (2023), and service demand in hospitals Tuominen et al. (2024); Bårdsen and Nymoen (2025). We also emphasize the importance of critically examining the potential risks involved in applying foundation models to econometric problems.
7 Declaration of Generative AI and AI-assisted technologies in the writing process
@@
8 CRediT author statement
Author 1: Data curation; Formal analysis; Investigation; Validation; Visualization; Writing- Original draft preparation. Author 2: Formal analysis; Investigation; Methodology; Validation; Visualization; Writing- Original draft preparation. Author 3: Conceptualization; Methodology; Project administration; Supervision; Validation; Visualization; Writing- Original draft preparation; Writing- Reviewing and Editing.

### v7

@@ 5.1 Actual over Expected Ratio
Table 2 reports the summary statistics over the out-of-sample period from January 2015 to September 2023, across 92 stocks, of the $|1-\text{AE}|$ values for the VaR forecasts at different levels, and the corresponding models in the out-sample period. Specifically, we report the minimum, mean, median, maximum, and standard deviation across all the stocks. Additionally, we report the count of stocks for which each of the considered models was the best (achieved lowest value of $|1-AE|$) or was within top two models (1st-2nd best). Note that, as mentioned before, the VaR forecasts for TimesFM pretrained model are only available at $10\%$, there we have three extra models in the last block of Table 2.
@@
Tables 4, 5 and 6 present the out-of-sample results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The out-of-sample results are also depicted via bar plots in Figures 2-4 in Appendix I. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. The higher the count, the better the model. From these tables, it is evident that the TimesFM shows promising results; in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or performs fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lengths show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@ 6 Conclusion and discussion
Determining whether pre-trained foundation time-series models can outperform traditional econometric methods is a key question, as their success would mean that anyone with minimal statistical and mathematical knowledge could achieve relatively good VaR forecasts. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training that yielded suboptimal results. However, fine-tuning the model with domain-specific data led to significant statistical improvements: In terms of Actual vs. Expected violations (AE), fine-tuned TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in the recent econometric literature [6]. Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as foundation models can be utilized with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect [36]. This lack of transparency is especially problematic in high-stakes financial settings, where understanding the reasoning behind predictions is essential for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) forecasting by offering a completely new alternative to traditional econometric methods. We anticipate future research addressing various econometric and financial market challenges, such as portfolio optimization and volatility forecasting, as well as applications in other fields where time series forecasting methods are required. These include forecasting electricity and commodity prices [37], macroeconomic variables [38], and service demand in hospitals [39]. We also emphasize the importance of critically examining the potential risks involved in applying foundation models to econometric problems.
Declaration of Generative AI and AI-assisted technologies in the writing process
@@
Acknowledgements
The first author acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie Actions (Grant Agreement No. 101150609, Project: HiddenTipChains).
Conflicts of Interest
The authors declare no conflicts of interest.
Appendix I